# Hard Examples Are All You Need: Maximizing GRPO Post-Training Under Annotation Budgets

### v2

@@ 5 Conclusion
This paper investigated a critical question for resource-constrained language model fine-tuning: under a fixed budget for example selection, which difficulty-based strategy maximizes the effectiveness of GRPO? Through comprehensive experiments across multiple models and tasks, we established that training on the hardest examples consistently yields superior performance on reasoning benchmarks. Deeper analysis showed that training on harder examples gives the model more opportunities to learn, and this gain generalizes to harder, out-of-distribution test sets.
@@ A.1 Reinforcement Learning from Human Feedback and GRPO
Reinforcement learning from human feedback (RLHF) has emerged as a dominant paradigm for aligning language models with human preferences [3, 17]. The standard RLHF pipeline involves training a reward model from human preferences, then optimizing the language model policy using algorithms like Proximal Policy Optimization (PPO) [19]. However, PPO-based RLHF suffers from high computational costs, training instability, and significant variance in gradient estimates.
Group Relative Policy Optimization (GRPO), introduced in DeepSeekMath [20], addresses these challenges by reformulating the optimization problem. Instead of training a separate reward model, GRPO directly optimizes using group-relative advantages computed from multiple samples per prompt. For each training prompt, GRPO samples a group of $G$ outputs, computes their rewards (e.g., correctness on math problems), and uses the deviation from the group mean as the advantage signal. This approach reduces variance through the group baseline while eliminating the need for reward model training.
@@ A.2 Example Difficulty and Data Selection
Curriculum learning suggests that training on examples in order of increasing difficulty can improve convergence and final performance [8]. However, curriculum learning typically assumes access to the full dataset and focuses on presentation order rather than subset selection. Our work differs by operating under a strict budget constraint where only a fraction of examples can be used.
Data valuation and influence methods aim to identify the most valuable training examples. Data Shapley [7] uses cooperative game theory to assign value to each training example based on its marginal contribution. However, these methods are computationally expensive and designed for supervised learning rather than RLHF settings.
Coreset selection techniques identify representative subsets that approximate the full dataset’s gradient [11, 10]. GradMatch selects examples whose gradients best match the full dataset gradient, while GLISTER focuses on generalization-based selection. These methods provide theoretical guarantees but require access to labels and gradient computation that may not be available in our budget-constrained setting.
Example forgetting and memorization studies have shown that neural networks exhibit consistent patterns in which examples are learned or forgotten during training [24].
Synthetic data is often an area of investigation for training, especially since it is often easier to control for difficulty. Our work doesn’t explicitly investigate using synthetic data and instead selects from a real, static larger dataset [25, 23, 9].
”Unforgettable” examples are learned early and retained, while others are repeatedly learned and forgotten. The Data Diet work [18] leverages Error L2-Norm (EL2N) scores computed early in training to identify important examples, showing that networks can be trained on small subsets without performance loss.
@@ A.3 Reasoning Benchmarks and Evaluation
GSM8K [4] has become the standard benchmark for grade-school mathematical reasoning, containing 8,792 problems requiring multi-step arithmetic and logical reasoning. The dataset’s problems are linguistically diverse but mathematically elementary, making it ideal for studying reasoning without confounding advanced mathematical knowledge.
BIG-Bench Hard (BBH) [22] curates 23 challenging tasks from the broader BIG-Bench suite where language models initially showed poor performance. The Tracking Shuffled Objects task, which we use in our experiments, requires models to track entity positions through a series of swapping operations: a pure reasoning task with minimal knowledge requirements.
Chain-of-thought prompting and self-consistency [26] have proven effective for improving reasoning performance. Self-consistency samples multiple reasoning paths and aggregates answers, often yielding substantial accuracy improvements. We leverage this insight in our difficulty estimation, using multi-sample success rates to robustly measure example hardness.
@@ A.4 GRPO
The GRPO algorithm operates by sampling groups of outputs for each training prompt and using relative performance within each group to compute advantages. Specifically, for each prompt $q$ in the selected subset, the algorithm samples $G$ outputs $\{o_{1},...,o_{G}\}$ from the current policy $\pi_{\theta}$, computes binary rewards $r_{i}\in\{0,1\}$ based on correctness, and calculates a group baseline $\bar{r}=\frac{1}{G}\sum_{i=1}^{G}r_{i}$. The advantage for each output is then computed as $A_{i}=r_{i}-\bar{r}$, which naturally centers the learning signal around the group’s average performance. These advantages drive policy updates while KL regularization to a reference policy $\pi_{\text{ref}}$ prevents catastrophic distribution shifts.
@@
$\mathcal{L}_{\text{GRPO}}=-\mathbb{E}_{q\sim S}\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}\left[A(o,q)\cdot\log\pi_{\theta}(o|q)-\beta\cdot\text{KL}(\pi_{\theta}||\pi_{\text{ref}})\right]$
@@ A.5 Closest GRPO Work
Difficulty-targeted selection and replay. Sun et al. propose a data-efficiency recipe for GRPO-style RL fine-tuning that (i) targets adaptive difficulty online, prioritizing questions of moderate difficulty expected to yield informative gradients, and (ii) introduces rollout replay to reuse recent samples and lower per-step cost. Across six LLM-dataset pairs they report 25-65% faster time-to-target performance than vanilla GRPO, suggesting that what you train on (and when) can matter as much as how much you train. Our paper shares the focus on difficulty-aware allocation under constraints, but differs in scope: we study budgeted, offline subset selection (hard vs. middle vs. easy vs. random) with a fixed labeling budget and analyze why ”hardest” can dominate on reasoning tasks, whereas Sun et al. optimize online time-to-quality with an adaptive, typically moderate-difficulty target and replay [21].
@@
Process rewards compatible with GRPO. Cui et al. (PRIME) tackle the sparsity and credit-assignment limits of outcome-only rewards by learning implicit process reward models online from outcome labels alone, then combining dense token-level signals with outcome rewards during RL. PRIME is compatible with standard advantage formulations (including GRPO) and removes a separate reward-modeling stage; empirically, they show sizable gains on competitive math and coding (e.g.,  15% average over SFT with a 7B base). Our study focuses on which examples to buy under an outcome-reward GRPO setup; PRIME is complementary, and our framework could be extended to ask whether difficulty-aware selection interacts with dense process rewards in similar ways (e.g., whether ”hardest” remains optimal when per-step feedback reduces variance) [5].
@@
GRPO biases and token efficiency. Liu et al. present a critical analysis of R1-Zero-like training and identify an optimization bias in GRPO: the objective can inflate response length, particularly for incorrect outputs, harming token efficiency. They introduce Dr. GRPO, an unbiased variant that preserves reasoning quality while improving efficiency, and demonstrate strong AIME-24 results with a 7B base model. For our budget-aware study, their findings motivate length-controlled evaluations and guardrails when comparing difficulty-conditioned subsets (e.g., ensuring that harder-subset gains are not confounded by pathological length growth and considering Dr. GRPO as a robustness check) [15].
@@
Positioning. Together, these works closest to GRPO underscore three levers we explicitly separate and test: (a) example selection by difficulty (Sun et al.), (b) reward shaping via dense process feedback (Cui et al.), and (c) objective design to avoid length bias (Liu et al.). Our contribution is orthogonal and complementary: a budget-aware, offline selection protocol and theory-informed analysis showing why, under outcome-level GRPO, training on the hardest decile can sustain a higher fraction of ”learnable” examples (non-degenerate group variance) and thus larger gains on reasoning tasks, while we adopt length controls in line with Dr. GRPO insights [21, 5, 15].
A.6 Out-Of-Distribution
Recent research has increasingly focused on evaluating large language models (LLMs) in out-of-distribution (OOD) scenarios, highlighting their potential and limitations in such contexts. Previous work conducted an empirical study on LLaMA models ranging from 7B to 65B parameters, demonstrating that simple cosine distance-based OOD detectors outperform traditional methods. Another work introduced LLM-GOOD, a framework that integrates LLMs’ zero-shot capabilities with graph neural networks to efficiently detect OOD nodes in text-attributed graphs, thereby reducing the need for extensive human annotation. Additionally, recent works have investigated the performance of reward models in OOD settings, as well as more traditional OOD evaluation [14, 12, 13, 27, 28]. These studies collectively underscore the evolving landscape of LLM evaluation in OOD contexts, emphasizing the need for tailored approaches to assess model reliability beyond standard benchmarks
A.7 Positioning Our Contributions
@@ Appendix B Hyperparameters
We set the group size $G=8$ for our main experiments based on preliminary studies showing this balances variance reduction with computational efficiency. The KL coefficient $\beta=0.1$ provides sufficient regularization without overly constraining learning. Training proceeds for 1000 steps with a learning rate of $3\times 10^{-5}$ using cosine decay, processing batches of 32 prompts per gradient update. We evaluate model performance every 100 steps to track learning dynamics, using the initial SFT checkpoint as our reference policy throughout training.

### v3

@@ 5 Conclusion
This paper investigated a critical question for resource-constrained language model fine-tuning: under a fixed budget for example selection, which difficulty-based strategy maximizes the effectiveness of GRPO? Through comprehensive experiments across multiple models and tasks, we established that training on the hardest examples consistently yields superior performance on reasoning benchmarks.
We validated this “hard examples are all you need” principle from two complementary perspectives. First, our difficulty-based selection experiments showed that examples with the lowest pass@k success rates drive performance gains up to 47%, while easy examples yield minimal improvements. Second, our analysis of training data categorized by base model performance revealed that nearly all GRPO learning value concentrates in examples the base model initially fails to solve. Training exclusively on examples where the base model achieves low success rates ($\hat{p}(x)<0.25$) yields up to 23.5% relative improvements compared to training on examples the base model already solves correctly ($\hat{p}(x)\geq 0.25$).
This convergence of evidence from both perspectives confirms a fundamental insight: GRPO learning requires training at the frontier of model capabilities. Examples the base model already masters reliably (high pass@k or those it solves correctly) are effectively the easiest and contribute minimal learning value, while examples where the model struggles (low pass@k or those it fails to solve) provide the variance in outcomes necessary for GRPO’s contrastive learning mechanism. This unified understanding provides practitioners with multiple practical approaches for identifying high-value training data while avoiding circular dependencies in dataset curation.
@@ A.1 Reinforcement Learning from Human Feedback and GRPO
Reinforcement learning from human feedback (RLHF) has emerged as a dominant paradigm for aligning language models with human preferences Christiano et al. (2017); Ouyang et al. (2022). The standard RLHF pipeline involves training a reward model from human preferences, then optimizing the language model policy using algorithms like Proximal Policy Optimization (PPO) Schulman et al. (2017). However, PPO-based RLHF suffers from high computational costs, training instability, and significant variance in gradient estimates.
Group Relative Policy Optimization (GRPO), introduced in DeepSeekMath Shao et al. (2024), addresses these challenges by reformulating the optimization problem. Instead of training a separate reward model, GRPO directly optimizes using group-relative advantages computed from multiple samples per prompt. For each training prompt, GRPO samples a group of $G$ outputs, computes their rewards (e.g., correctness on math problems), and uses the deviation from the group mean as the advantage signal. This approach reduces variance through the group baseline while eliminating the need for reward model training.
@@ A.2 Example Difficulty and Data Selection
Curriculum learning suggests that training on examples in order of increasing difficulty can improve convergence and final performance Hacohen & Weinshall (2019). However, curriculum learning typically assumes access to the full dataset and focuses on presentation order rather than subset selection. Our work differs by operating under a strict budget constraint where only a fraction of examples can be used.
Data valuation and influence methods aim to identify the most valuable training examples. Data Shapley Ghorbani & Zou (2019) uses cooperative game theory to assign value to each training example based on its marginal contribution. However, these methods are computationally expensive and designed for supervised learning rather than RLHF settings.
Coreset selection techniques identify representative subsets that approximate the full dataset’s gradient Killamsetty et al. (2021a; b). GradMatch selects examples whose gradients best match the full dataset gradient, while GLISTER focuses on generalization-based selection. These methods provide theoretical guarantees but require access to labels and gradient computation that may not be available in our budget-constrained setting.
Example forgetting and memorization studies have shown that neural networks exhibit consistent patterns in which examples are learned or forgotten during training Toneva et al. (2019).
Synthetic data is often an area of investigation for training, especially since it is often easier to control for difficulty. Our work doesn’t explicitly investigate using synthetic data and instead selects from a real, static larger dataset Trabucco et al. (2025); Tian et al. (2025); Havrilla et al. (2025).
“Unforgettable” examples are learned early and retained, while others are repeatedly learned and forgotten. The Data Diet work Paul et al. (2021) leverages Error L2-Norm (EL2N) scores computed early in training to identify important examples, showing that networks can be trained on small subsets without performance loss.
@@ A.3 Reasoning Benchmarks and Evaluation
GSM8K Cobbe et al. (2021) has become the standard benchmark for grade-school mathematical reasoning, containing 8,792 problems requiring multi-step arithmetic and logical reasoning. The dataset’s problems are linguistically diverse but mathematically elementary, making it ideal for studying reasoning without confounding advanced mathematical knowledge.
BIG-Bench Hard (BBH) Suzgun et al. (2023) curates 23 challenging tasks from the broader BIG-Bench suite where language models initially showed poor performance. The Tracking Shuffled Objects task, which we use in our experiments, requires models to track entity positions through a series of swapping operations: a pure reasoning task with minimal knowledge requirements.
Chain-of-thought prompting and self-consistency Wang et al. (2022) have proven effective for improving reasoning performance. Self-consistency samples multiple reasoning paths and aggregates answers, often yielding substantial accuracy improvements. We leverage this insight in our difficulty estimation, using multi-sample success rates to robustly measure example hardness.
@@ A.4 GRPO
[…]
@@
[…]
@@ A.5 Closest GRPO Work
Difficulty-targeted selection and replay. Sun et al. propose a data-efficiency recipe for GRPO-style RL fine-tuning that (i) targets adaptive difficulty online, prioritizing questions of moderate difficulty expected to yield informative gradients, and (ii) introduces rollout replay to reuse recent samples and lower per-step cost. Across six LLM-dataset pairs they report 25-65% faster time-to-target performance than vanilla GRPO, suggesting that what you train on (and when) can matter as much as how much you train. Our paper shares the focus on difficulty-aware allocation under constraints, but differs in scope: we study budgeted, offline subset selection (hard vs. medium vs. easy vs. random) with a fixed labeling budget and analyze why “hardest” can dominate on reasoning tasks, whereas Sun et al. optimize online time-to-quality with an adaptive, typically moderate-difficulty target and replay Sun et al. (2025).
@@
Process rewards compatible with GRPO. Cui et al. (PRIME) tackle the sparsity and credit-assignment limits of outcome-only rewards by learning implicit process reward models online from outcome labels alone, then combining dense token-level signals with outcome rewards during RL. PRIME is compatible with standard advantage formulations (including GRPO) and removes a separate reward-modeling stage; empirically, they show sizable gains on competitive math and coding (e.g.,  15% average over SFT with a 7B base). Our study focuses on which examples to buy under an outcome-reward GRPO setup; PRIME is complementary, and our framework could be extended to ask whether difficulty-aware selection interacts with dense process rewards in similar ways (e.g., whether “hardest” remains optimal when per-step feedback reduces variance) Cui et al. (2025).
@@
GRPO biases and token efficiency. Liu et al. present a critical analysis of R1-Zero-like training and identify an optimization bias in GRPO: the objective can inflate response length, particularly for incorrect outputs, harming token efficiency. They introduce Dr. GRPO, an unbiased variant that preserves reasoning quality while improving efficiency, and demonstrate strong AIME-24 results with a 7B base model. For our budget-aware study, their findings motivate length-controlled evaluations and guardrails when comparing difficulty-conditioned subsets (e.g., ensuring that harder-subset gains are not confounded by pathological length growth and considering Dr. GRPO as a robustness check) Liu et al. (2025).
@@
Positioning. Together, these works closest to GRPO underscore three levers we explicitly separate and test: (a) example selection by difficulty (Sun et al.), (b) reward shaping via dense process feedback (Cui et al.), and (c) objective design to avoid length bias (Liu et al.). Our contribution is orthogonal and complementary: a budget-aware, offline selection protocol and theory-informed analysis showing why, under outcome-level GRPO, training on the hardest decile can sustain a higher fraction of “learnable” examples (non-degenerate group variance) and thus larger gains on reasoning tasks, while we adopt length controls in line with Dr. GRPO insights Sun et al. (2025); Cui et al. (2025); Liu et al. (2025).
A.6 Positioning Our Contributions
@@ Appendix B Hyperparameters
We set the group size $G=8$ for our main experiments based on preliminary studies showing this balances variance reduction with computational efficiency. The KL coefficient $\beta=0.1$ provides sufficient regularization without overly constraining learning. Training proceeds for 1000 steps with a learning rate of $3\times 10^{-5}$ using cosine decay, processing 1 prompt per gradient update due to memory constraints. We use 4-bit quantized models from Unsloth Daniel Han & team (2023). We evaluate model performance every 100 steps to track learning dynamics, using the initial SFT checkpoint as our reference policy throughout training.
Appendix C LLM Use
Large Language Models (LLMs) were used for the following purposes during this research:
  1. 1.
    Writing some of the code for the experiments: LLMs assisted in generating and debugging portions of the experimental code, including data processing scripts and evaluation utilities.
  2. 2.
    Proofreading the text in the paper: LLMs were used to identify grammatical errors, improve clarity, and ensure consistency in the manuscript.
  3. 3.
    LaTeX support: LLMs provided assistance with LaTeX syntax for complex table formatting and figure placement.